# Statistical features of quantum chaos using the Krylov operator complexity

### v1

@@ 4 Statistical properties of Lanczos coefficients
In the other aspect, suppose the variance $\sigma_{\mathcal{O}}^{2}$ is computed from $2K$ Lanczos coefficients of a given initial operator $\mathcal{O}^{(0)}$, with $2K$ being the algorithm-steps selected from the Krylov space. By definition (2.7), we have
$\sigma_{\mathcal{O}}^{2}$ $=\frac{1}{K}\sum_{i=1}^{K}\bigg{(}x_{i}(\mathcal{O}^{(0)};E_{mn})\bigg{)}^{2}-\frac{1}{K^{2}}\bigg{(}\sum_{i=1}^{K}x_{i}(\mathcal{O}^{(0)};E_{mn})\bigg{)}^{2}$
$=(\frac{1}{K}-\frac{1}{K^{2}})\sum_{i=1}^{K}\bigg{(}x_{i}(\mathcal{O}^{(0)};E_{mn})\bigg{)}^{2}-\frac{1}{K^{2}}\sum_{i\neq j}^{K}x_{i}(\mathcal{O}^{(0)};E_{mn})x_{j}(\mathcal{O}^{(0)};E_{mn}).$ (4.9)
[…]
$\sum_{i\neq j}^{K}\langle x_{i}(\mathcal{O}^{(0)};E_{mn})x_{j}(\mathcal{O}^{(0)};E_{mn})\rangle\sim\sum_{i\neq j}^{K}\langle x_{i}(\mathcal{O}^{(0)};E_{mn})\rangle\langle x_{j}(\mathcal{O}^{(0)};E_{mn})\rangle\sim 0.$ (4.10)
This qualitatively can be understood from an analog with random walks and from the nature of the initial distribution (GOE, GUE and the uniform distribution). So after taking the average over initial random operators, we obtain qualitatively
$\langle\sigma_{\mathcal{O}}^{2}\rangle$ $\sim(\frac{1}{K}-\frac{1}{K^{2}})\sum_{i=1}^{K}\bigg{\langle}x_{i}(\mathcal{O}^{(0)};E_{mn})^{2}\bigg{\rangle}.$ (4.11)
If the sample is large enough, we can qualitatively expect $\sigma_{\mathcal{O}}^{2}$ to follow the normal distribution as a result of the central limit theorem.
@@
From the above qualitative analysis, we see that the numbers of random numbers are determined by $\mathcal{N}_{max}$, $K$ and sample size of initial operators. If they are large enough, we would expect the resulting distribution to be normal. In the following computation, we use the choice $\mathcal{N}_{max}=100$ and $K=500$ in consistent with Hashimoto:2023swv , and we average over $5000$ intial operators. This number turns out to be large enough and the resulting distribution function $f_{\sigma^{2}}$ is the normal distribution. (In Appendix C, we check the case $\mathcal{N}_{max}=5$. The resulting distribution $f_{\sigma^{2}}$ is not normal there, and interesting, the phenomenon of separation of $f_{\sigma^{2}}$ for different initial distributions disappears.)
@@ 5 Model and Results
The region $\Omega$ is obtained by cutting out a circle of radius $l$ from an equilateral triangle of side length $L$, as shown in Fig.1. The system is integrable when $l=0$ and chaotic when $0<l<L$. For convenient, the area of $\Omega$ is kept at 1 and the parameter $a\equiv l/L$ is introduced, whose value determines the degree of chaos for the system.
@@
For practical computation in the energy representation, firstly we set the value $\mathcal{N}_{max}$ to truncate the maximum number of energy levels. Here we adopt the convention $\mathcal{N}_{max}=100$ in accordance with Hashimoto:2023swv . Secondly, we pick a random matrix $\mathcal{M}_{\mathcal{O}}$ of size $100\times 100$ from the probability distributions (3.2) and (3.3), which is the matrix of the initial random operator $\mathcal{O}$. Then we compute the Lanczos coefficient $\left\{b_{n}\right\}$ of $\mathcal{K}_{m}(\mathcal{L},\mathcal{M}_{\mathcal{O}})$ by the algorithm in Section 2. To calculate the variance $\sigma^{2}$ of the Lanczos coefficients, we choose $5\mathcal{N}_{max}$ coefficients in consistent with Hashimoto:2023swv . The set of coefficients $\left\{b_{n}\right\}$ is from the algorithm step sizes $500\leq n\leq 1000$, as shown in the Fig.2(a). Note that this selection of the set of $\left\{b_{n}\right\}$ does not qualitatively affect the result of this paper. The effect of selecting the set of $\left\{b_{n}\right\}$ from different algorithm step sizes is discussed in the appendix B. Finally, we repeat this process $5000$ times to get the resulting distribution function $f(\mathcal{O})$.
@@
For the different distributions (GOE, GUE, URE, UIM and UCP) of initial operators, the sampled variances $\sigma^{2}$ are shown in Fig. 3, where $m$ labels the samples. For visual cleanness, we only show $1000$ samples instead of all $5000$ samples. For GOE and GUE, they overlap completely in the nonchaotic case in Fig. 3(b), and separate from each other in the chaotic case in Fig. 3(a). Similary for the uniform distributions, they mix together in the nonchaotic case in Fig. 3(d), and split into two separated groups in the chaotic case in Fig. 3(c), with URE and UIM being one group and UCP being the other.
@@
The corresponding histogram of the sampled $\sigma^{2}$ is shown in Fig. 4, which shows that the resulting probability distribution resembles the normal distribution. For the Gaussian ensembles, we can see clearly that they are overlapped in the nonchaotic case in Fig. 4(b) and separated in the chaotic case in Fig. 4(a). For the uniform distributions, they are overlapped in the nonchaotic case in Fig. 4(d). In the chaotic case in Fig. 4(c), URE and UIM are still together, but they are separated from UCP.
@@
Figure 4: The histogram of variances $\sigma^{2}$ after sampling the initial operator 5000 times from GOE, GUE, URE, UIM and UCP. Obviously they resemble normal distributions.
We have shown the Gaussian ensembles and the uniform distributions separately for visual cleanness in the above figures 3 and  4. Their axis values indicate that they can be compared together in one figure. For this purpose, we fit 111 The details of this data fitting is shown in the Appendix A. their histogram to the normal distribution and find the average value $\langle\sigma^{2}\rangle$ and the standard deviation, which characterize the resulting distribution $f_{\sigma^{2}}$. Then the resulting distribution $f_{\sigma^{2}}$ can be shown in a figure where the point is this average value and the error bar is the standard deviation. For all different distributions (GOE, GUE, URE, UIM and UCP) of initial operators, the resulting distributions $f_{\sigma^{2}}$ are shown in Fig. 5, as the system varies from nonchaotic $a=0$ to chaotic $a=1$. We see that when the system is nonchaotic, the data points of all distributions overlap together within the range of error bars. As the system becomes chaotic, different distributions start to separate at $a=0.1$ and the distance of separation is roughly fixed after $a=0.15$. Note that the axis is logarithmic, so the separation is very large. We see that they split into two groups well-separated: one group is of GOE, URE and UIM, the other group is GUE and UCP.
@@
Figure 5: The logarithm plot of the average $\langle\sigma^{2}\rangle$ of resulting normal distributions $f_{\sigma^{2}}$, for all different distributions (GOE, GUE, URE, UIM and UCP) of initial operators. The error bar is the standard deviation. The system dynamics changes from nonchaotic to chaotic, as $a$ varies from 00 to $1$.
6 Discussion
Firstly, we emphasize the importance of the resulting distribution being the normal distribution. We repeat the computation for $\mathcal{N}_{max}=5$ in Appendix C, where the resulting distribution $f_{\sigma^{2}}$ is not normal and there is no separation of $f_{\sigma^{2}}$ in the chaotic case. This suggests that the resulting distributions being normal is essential for their separation behavior in the chaotic case.
Qualitatively we can understand this by the central limit theorem as discussed in Section 4. The resulting distribution being normal means that we have enough random numbers in the sample, so that the statistics is reliable and robust. Luckily, the typical choice of $\mathcal{N}_{max}=100$ in literature is adequate for the normal distribution. It remains an open problem to quantitatively understand the connection between being normal distributions and the separation behavior in the chaotic case.
Secondly, we emphasize that these overlapping-separating behaviors do not depend on the choice of algorithm step-sizes of coefficients $\left\{b_{n}\right\}$. In appendix B, we choose the step-sizes $1000\leq n\leq 1500$ from the Krylov space in Fig. 2. We still get the resulting normal distributions and the overlapping-splitting behaviors. Note that the detailed values of the average and the standard deviation will change, it is the overlapping-separating behavior that does not change. It remains an open problem to quantitatively prove this independence on different choices of algorithm step-sizes.
Thirdly, we discuss the resulting two groups of separation. The GOE, URE and UIM are in one group. We can view GOE, URE and UIM as essentially composed of only real numbers, because the UIM is just real numbers multiplied by the overall factor-the imaginary $i$. The other group is GUE and UCP, which are essentially composed of complex numbers with both real and imaginary parts. So one group is essentially related with one-dimensional random numbers and the other group is related with two-dimensional random numbers. This distinction between real and complex numbers may be related to the nature of the integrability-breaking-term, or to the symmetry of the model. This remains to be one of the open problems.
Finally, for the chaotic case, we compare our separation behavior with a similar behavior-of-separation in RMT Rabinovici:2022beu . For a spin system Rabinovici:2022beu under chaotic dynamics, the Hamiltonian is sampled from RMT ensembles while the initial operator is fixed. For Hamiltonians from GUE and GOE, a separation is obtained in the late-time saturation of the Krylov complexity. Qualitatively, the behavior of late-time saturation of K-complexity are consistent with the behavior of variances $\sigma^{2}$ of the Lanczos coefficients Rabinovici:2021qqt . So it might be possible that the ensemble average of the variance $\sigma^{2}$ over RMT Hamiltonian are also separating in Rabinovici:2022beu , although it is not computed there. If this indeed happen, the separation behavior of the resulting distributions $f_{\sigma^{2}}$ and of the late-time saturation of the Krylov complexity can be viewed as ’dual’ to each other for chaotic systems.
Qualitatively, this ’duality’ can be understood as following: in one side, the Hamiltonian are random and the initial operator is fixed, and in the other side, the Hamiltonian is fixed and the initial operators are random. They are ’equivalent’ when taking the statistics, at least for chaotic systems. From the proof (4.7) of the lemma in Section 4, we see that the Lanczos coefficients and the Krylov basis depend only on the initial operators and energy levels. Both the random energy levels and random initial operators can give random Lanczos coefficients, so intuitively, these statistics can be viewed as dual or equivalent. It remains an open problem to quantitatively prove this connection.
In conclusion, we pick the Sinai billiard and generate random initial operators from given probability distribtions (GOE, GUE, URE, UIM and UCP). Then we find that the resulting distributions of the variance of Lanzcos coefficients are normal distributions. They are overlapped in the nonchaotic case. When the system becomes chaotic, they split into two groups, with one group composed of GOE, URE and UIM, and the other group composed of GUE and UCP. We analyze the reason of being normal and discuss the overlap-separation behavior with open problems.

### v2

@@ 3.3 Distribution $f_{\sigma^{2}}$ of the variance $\sigma^{2}$
For convenience, suppose the variance $\sigma_{\mathcal{O}}^{2}$ is computed from $K=2p$ Lanczos coefficients for a given initial operator $\mathcal{O}$. By definition (3.3), we have
$\sigma_{\mathcal{O}}^{2}$ $=\frac{1}{p}\sum_{i=1}^{p}x_{i}^{2}-\frac{1}{p^{2}}\bigg{(}\sum_{i=1}^{p}x_{i}\bigg{)}^{2}=\bigg{(}\frac{1}{p}-\frac{1}{p^{2}}\bigg{)}\sum_{i=1}^{p}x_{i}^{2}-\frac{1}{p^{2}}\sum_{i\neq j}^{p}x_{i}x_{j}.$ (3.17)
Qualitatively, we can expect that the second term with $i\neq j$ can be neglected compared with the first term. This is true when the set size $K$ is large. In the detailed computation of this paper, this condition is satisfied and the second term is smaller than the first term by at least one order of magnitude. So we obtain qualitatively
$\sigma_{\mathcal{O}}^{2}$ $\sim\bigg{(}\frac{1}{p}-\frac{1}{p^{2}}\bigg{)}\sum_{i=1}^{p}x_{i}^{2}.$ (3.18)
In the previous subsection 3.2, we have discussed the physics that $x_{i}$ are independent random variables from a univariate normal distribution $N(0,\sigma_{x})$. If the standard deviation $\sigma_{x}$ is known, we can scale $x_{i}$ by an overall factor to make its standard deviation being $1$. In this way, the resulting distribution $f_{\sigma^{2}}$ (3.18) is proportional to the chi-square distribution by an overall factor, called the rescaled chi-square distribution here. When $\mathcal{N}_{max}$ increases, it would approach the normal distribution.
Numerical results indeed support this expectation. In Figure 2, we show the histogram of the variance $\sigma^{2}$ for the case of GOE when the system is chaotic. When $\mathcal{N}_{max}=5$, the histogram clearly satisfies the rescaled chi-square distribution with the curve being the fitted distribution. As it increases to $\mathcal{N}_{max}=100$, the histogram approaches the normal distribution and the curve are the fitted normal distribution. Similar properties are there for the other ensembles and for the nonchaotic case, as discussed in Section 4 and Appendix C.
(a) $\mathcal{N}_{max}=5$, rescaled chi-square distribution.
(b) $\mathcal{N}_{max}=100$, normal distribution. Skewness=0.201339, Kurtosis=3.19744
Figure 2: Chi-square distribution of $f_{\sigma^{2}}$ in the chaotic case for GOE.
@@
From this section’s analysis, we see that the amount of random numbers is determined by $\mathcal{N}_{max}$, $K$ and the sample size $N$ of initial operators. If they are large enough, we would expect the resulting distribution to be normal. In the following computation, we use the choice $\mathcal{N}_{max}=100$ and $K=500$ in consistent with Hashimoto:2023swv , and we average over $N=5000$ intial operators. This number turns out to be large enough and the resulting distribution function $f_{\sigma^{2}}$ is the normal distribution.
@@ 4 Model and Results
The region $\Omega$ is obtained by cutting out a circle of radius $l$ from an equilateral triangle of side length $L$, as shown in Figure 3. The system is integrable when $l=0$ and chaotic when $0<l<L$. For convenient, the area of $\Omega$ is kept at 1 and the parameter $a\equiv l/L$ is introduced, whose value determines the degree of chaos for the system.
@@
For practical computation in the energy representation, we need to set the value of $\mathcal{N}_{max}$ to truncate the number of energy levels. It is adequate to take a value as long as the resulting distributions of $\left\langle x_{i}x_{j}\right\rangle$ and $f_{\sigma^{2}}$ are normal. When this happens, there are enough randomness in the statistics so that $\left\langle x_{i}x_{j}\right\rangle$ and $f_{\sigma^{2}}$ can distinguish the chaotic and the nonchaotic behaviors. Empirically, $\mathcal{N}_{max}=50$ is enough here. For the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$, we use $\mathcal{N}_{max}=50$ for the purpose of numerical efficiency. For the distribution $f_{\sigma^{2}}$ of the variance, we use $\mathcal{N}_{max}=100$ to keep accordance with Hashimoto:2023swv . Then we pick a random matrix $\mathcal{M}_{\mathcal{O}}$ of size $\mathcal{N}_{max}\times\mathcal{N}_{max}$ from the probability distributions (2.8)-(2.10) for the initial random operator $\mathcal{O}$. After that, we compute the Lanczos coefficient $\left\{b_{n}\right\}$ of $\mathcal{K}_{m}(\mathcal{L},\mathcal{M}_{\mathcal{O}})$ by the algorithm in Section 2. To calculate the variance $\sigma^{2}$ of Lanczos coefficients, we choose the set size to be $K=5\mathcal{N}_{max}$, in consistent with Hashimoto:2023swv . Finally, we repeat this process $N=5000$ times to get the resulting distribution of $\left\langle x_{i}x_{j}\right\rangle$ and $f_{\sigma^{2}}$.
In Figure 4, an example of Lanczos coefficients $b_{n}$ for the case of GUE is shown. There is a clear difference between the chaotic and the nonchaotic case, which can be explored further by the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$. The set of coefficients $\left\{b_{n}\right\}$ between the two red lines, $5\mathcal{N}_{max}\leq n\leq 10\mathcal{N}_{max}$, is selected to compute the variance $\sigma^{2}$. Note that this selection of the set does not qualitatively affect the result of this paper. The effect of selecting the set of $\left\{b_{n}\right\}$ from different algorithm step sizes is discussed in the appendix B.
In Figure 5, the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$ is shown for the case of GUE. With $\mathcal{N}_{max}=50$, its histogram approaches the normal distribution as shown in Figure 1(b) so that there is enough randomness in the statistics. There is a clear distinction in the pattern between the chaotic and the nonchaotic case. In the nonchaotic case, there are heavy cross-like structures with large matrix values. But in the chaotic case, these heavy cross-like structure disappears (or fades away) and matrix values are more uniformly or smoothly distributed.
In Figure 6, the variances $\sigma^{2}$ are shown for the samples of different distributions (GOE, GUE, URE, UIM and UCP) of initial operators, where $m$ labels the sample. Note that only the first 1000 out of the total 5000 samples are shown for the purpose of visual clearness, and the remaining samples have similar behaviors. For GOE and GUE, they overlap completely in the nonchaotic case in Figure 6(b), and separate from each other in the chaotic case in Figure 6(a). Similary for the uniform distributions, they mix together in the nonchaotic case in Figure 6(d), and split into two separated groups in the chaotic case in Figure 6(c), with URE and UIM being one group and UCP being the other.
In Figure 7, the histogram of all the sampled $\sigma^{2}$ is shown and the resulting probability distribution resembles the normal distribution. For the Gaussian ensembles, we can see clearly that they are overlapped in the nonchaotic case in Figure 7(b) and separated in the chaotic case in Figure 7(a). For the uniform distributions, they are overlapped in the nonchaotic case in Figure 7(d). In the chaotic case in Figure 7(c), URE and UIM are still together, but they are separated from UCP.
Figures 6 and  7 straightforwardly show that the resulting distributions $f_{\sigma^{2}}$ can distinguish the chaotic and the nonchaotic behaviors. Their axis values indicate that they can be compared together in one plot. So we fit 333 The details of this data fitting is shown in the Appendix A. their histogram to the normal distribution and find the average value $\langle\sigma^{2}\rangle$ and the standard deviation, which can characterize the resulting distribution $f_{\sigma^{2}}$. Then the resulting distribution $f_{\sigma^{2}}$ can be shown in one plot where the point is this average value and the error bar is the standard deviation. In Figure 8, the resulting distributions $f_{\sigma^{2}}$ are shown for all different distributions (GOE, GUE, URE, UIM and UCP) of initial operators, as the system varies from nonchaotic $a=0$ to chaotic $a=1$. When the system is nonchaotic, the data points of all distributions overlap together within the range of error bars. As the system becomes chaotic, different distributions start to separate approximately at $a=0.1$ and the distance of separation is roughly fixed after $a=0.15$. Note that the axis is logarithmic, so the separation is very large. We see that they split into two groups well-separated: one group is of GOE, URE and UIM, the other group is GUE and UCP.
5 Discussion
To begin, the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$ has different patterns between the nonchaotic and the chaotic case, as shown in Figure 5. From the aspect of RMT, the average correlation matrix $\left\langle x_{i}x_{j}\right\rangle$ is the common estimator of the statistics. So different patterns qualitatively tell us that the statistics of $\left\langle x_{i}x_{j}\right\rangle$ can distinguish between the nonchaotic and the chaotic case. On the other hand, $x_{i}$ characterizes the disorderedness of Lanczos coefficients from the aspect of the Krylov chain. So $\left\langle x_{i}x_{j}\right\rangle$ quantitatively captures the erratic behavior and the correlation of Lanczos coefficients. In the nonchaotic case of Figure 5(b), the large matrix values indicate more erratic behavior of Lanczos coefficients and more localization on the Krylov chain. So it provides another view for the connection between chaos and Anderson localization and is consistent with results of Dymarsky:2019elm ; Rabinovici:2021qqt .
The repetition of the heavy cross-like structure in Figure 5(b) is also a very interesting phenomenon. It might come from some abstract periodic boundary conditions of the Krylov chain. Or it might be connected to the phenomenon of two-electron correlations of Anderson localization on the Krylov chain Weinmann1995h2eOF . In Figure 9, we plot the average correlation matrix of the wave function $\langle\ln{|\psi_{2m}\psi_{2n}|}\rangle$ for the zero frequency state of the Krylov chain. There are also heavy cross-like structures for the nonchaotic case in Figure 9(b). These cross-like structures locally look like the cross-like structure of the two-electron correlations of Anderson localization in disordered mesoscopic ring Weinmann1995h2eOF . And the cross-like structure of $\ln{|\psi_{2m}\psi_{2n}|}$ disappears in the chaotic case in Figure 9(a). The physics of this structure remains an open problem at this stage.
Then, we emphasize that these overlapping-separating behaviors do not depend on the choice of algorithm step-sizes of coefficients $\left\{b_{n}\right\}$. In appendix B, we choose the step-sizes $10\mathcal{N}_{max}\leq n\leq 15\mathcal{N}_{max}$ from the Krylov space in Figure 4. We still get the resulting normal distributions and the overlapping-splitting behaviors. Note that the detailed values of the average and the standard deviation will change, it is the overlapping-separating behavior that does not change. It remains an open problem to quantitatively prove this independence on different choices of algorithm step-sizes.
In addition, we emphasize the importance of the resulting distribution being the normal distribution. We repeat the computation for $\mathcal{N}_{max}=5$ in Appendix C, where the resulting distribution $f_{\sigma^{2}}$ is not normal and there is no separation of $f_{\sigma^{2}}$ in the chaotic case. This suggests that the resulting distributions being normal is essential for their separation behavior in the chaotic case.
Qualitatively we can understand this by the central limit theorem as discussed in Section 3. The resulting distribution being normal means that we have enough randomness in the sample, so that the statistics is reliable and robust. Empirically, a choice of $\mathcal{N}_{max}=50$ is already adequate to get the normal distribution. It remains an open problem to quantitatively understand the connection between being normal distributions and the separation behavior of $f_{\sigma^{2}}$.
After that, we discuss the resulting two groups of separation. The GOE, URE and UIM are in one group. We can view GOE, URE and UIM as essentially composed of only real numbers, because the UIM is just real numbers multiplied by the overall factor-the imaginary $i$. The other group is GUE and UCP, which are essentially composed of complex numbers with both real and imaginary parts. So one group is essentially related with one-dimensional random numbers and the other group is related with two-dimensional random numbers. This distinction between real and complex numbers may be related to the nature of the integrability-breaking-term, or to the symmetry of the model haake1991quantum . This remains to be one of the open problems.
Finally, we explain a connection (or equivalence) between the average over initial operators and the RMT Hamiltonians. For the chaotic case, a similar behavior-of-separation is obtained related with RMT Rabinovici:2022beu . For a spin system Rabinovici:2022beu under chaotic dynamics, the Hamiltonian is sampled from RMT ensembles while the initial operator is fixed. For Hamiltonians from GUE and GOE, a separation is obtained in the late-time saturation of the Krylov complexity. Qualitatively, the behavior of late-time saturation of K-complexity are consistent with the behavior of variances $\sigma^{2}$ of the Lanczos coefficients Rabinovici:2021qqt . So it might be possible that the ensemble average of the variance $\sigma^{2}$ over RMT Hamiltonian are also separating in Rabinovici:2022beu , although it is not computed there. If this indeed happen, the separation behavior of the resulting distributions $f_{\sigma^{2}}$ and of the late-time saturation of the Krylov complexity can be viewed as ’dual’ to each other for chaotic systems.
Another sign for this ’duality’ also appears in the study of the spin chain Trigueros:2021rwj . There the randomness is introduced by the random transverse field from a uniform distribution. So it is equivalent to a RMT Hamiltonian. There the statistics is focused on the parameters of the linear extrapolation formula of Lanczos coefficients. These fitted parameters also follow the normal distribution, similar to the resulting distribution $f_{\sigma^{2}}$, although they are completely different quantities. This suggests that the statistics would be normal as long as there is enough randomness, whether it is due to randomn initial operators or RMT Hamiltonians.
Qualitatively, this ’duality’ can be understood as follows: in one side, the Hamiltonian are random and the initial operator is fixed, and in the other side, the Hamiltonian is fixed and the initial operators are random. They are ’equivalent’ when taking the statistics, at least for chaotic systems. From the proof (3.10) of the lemma in Section 3, we see that the Lanczos coefficients and the Krylov basis depend only on the initial operators and energy levels. Both the random energy levels and random initial operators can give random Lanczos coefficients, so intuitively, these statistics can be viewed as dual or equivalent. It remains an open problem to quantitatively prove this duality. In Rabinovici:2021qqt , the Lanczos coefficients are rewritten in terms of the Hankel determinants of level spacings, which might be helpful in this direction.
In conclusion, we find two statistical quantities that have distinct behaviors as the Sinai billiard system changes from nonchaotic to chaotic, i.e., the average correlation matrix $\langle x_{i}x_{j}\rangle$ and the resulting distribution $f_{\sigma^{2}}$ of the variance of Lanczos coefficients. The statistics of these two quantities are the Wishart distribution and the rescaled chi-square distribution respectively, independent of distributions of initial operators. When the chi-square distribution approaches the normal distribution, the resulting distributions $f_{\sigma^{2}}$ can distinguish the nonchaotic and the chaotic regime: they are overlapped in the nonchaotic case and they split into two groups in the chaotic case. The average correlation matrix $\langle x_{i}x_{j}\rangle$ also exhibits different patterns in the nonchaotic and the chaotic regime: there are heavy cross-like structures in the nonchaotic case. These results provide a dual view on results in the literature and reveal a deeper connection among Lanczos coefficients, RMT and the Anderson localization.
@@ 5 Discussion
(a) $\left|\langle x_{i}x_{j}\rangle\right|$ in chaotic case $a=1$.
(b) $\left|\langle x_{i}x_{j}\rangle\right|$ in nonchaotic case $a=0$.
Figure 5: Patterns of $\left|\langle x_{i}x_{j}\rangle\right|$ with $\mathcal{N}_{max}=50$
@@
Figure 7: The histogram of variances $\sigma^{2}$ after sampling the initial operator 5000 times from GOE, GUE, URE, UIM and UCP. The property of being normal and the overlapping-separating behavior are obvious.
[t] Refer to caption
@@
Figure 8: The logarithm plot for the average $\langle\sigma^{2}\rangle$ of the resulting normal distribution $f_{\sigma^{2}}$, for all different distributions of initial operators. The error bar is the standard deviation. The system dynamics changes from nonchaotic to chaotic, as $a$ varies from 00 to $1$.
(a) $\left\langle\ln{|\psi_{2m}\psi_{2n}|}\right\rangle$ in chaotic case $a=1$.
(b) $\left\langle\ln{|\psi_{2m}\psi_{2n}|}\right\rangle$ in nonchaotic case $a=0$.
Figure 9: Patterns of $\left\langle\ln{|\psi_{2m}\psi_{2n}|}\right\rangle$ with $\mathcal{N}_{max}=50$